OpenStudy (anonymous):

Solve the Proportion x-4/6=x/8 a)2 b)16 c)12 d) 8

OpenStudy (anonymous):

.

OpenStudy (anonymous):

1. x/4 = 8/6 6x = 32 x = 5.33 or 5 1/3

OpenStudy (anonymous):

if it were one of the optionz, id give u a medal. !

OpenStudy (anonymous):

@Luis_Rivera do you know how to solve this?

OpenStudy (anonymous):